# Object-level Geometric Structure Preserving for Natural Image Stitching

### v1

@@ Abstract
The topic of stitching images with globally natural structures holds paramount significance. Current methodologies exhibit the ability to preserve local geometric structures, yet fall short in maintaining relationships between these geometric structures. In this paper, we endeavor to safeguard the overall, OBJect-level structures within images based on Global Similarity Prior, while concurrently mitigating distortion and ghosting artifacts with OBJ-GSP. Our approach leverages the Segment Anything Model [25] to extract geometric structures with semantic information, enhancing the algorithm’s ability to preserve objects in a manner that aligns more intuitively with human perception. We seek to identify spatial constraints that govern the relationships between various geometric boundaries. Recognizing that multiple geometric boundaries collectively define complete objects, we employ triangular meshes to safeguard not only individual geometric structures but also the overall shapes of objects within the images. Empirical evaluations across multiple image stitching datasets demonstrate that our method establishes a new state-of-the-art benchmark in image stitching. Our implementation and dataset is publicly available at https://github.com/RussRobin/OBJ-GSP
@@ 1 Introduction
Figure 1: Our panorama (d) preserves structures well even if the structure marked with red line is not recognized by SAM (b). By preserving the object-level structure of the building, the local shapes are also preserved.
Image stitching aims to align multiple images and create a composite image with a larger field of view. This method is widely utilized across diverse domains, including smartphone panoramic photography [43], robotic navigation [10], and virtual reality experiences [1, 24]. The field of image stitching algorithms grapples with two fundamental challenges: alignment and distortion. Autostitch [3] initially undertakes the extraction of feature points and computes the homography matrix to align feature points between images. Approaches such as APAP [47] and GSP [7] often segment the target image into smaller grids, computing local homographic transformations between the target and reference images within each grid. This, combined with global homography transformations, effectively addresses the alignment concern. However, a relentless pursuit of feature point alignment can inadvertently introduce distortion into the resulting images. In light of this, GES-GSP [13] commences with edge detection, extracting linear and curvilinear structures within the image. Subsequently, triangle sampling is applied to the curves. During the grid deformation process, consideration is given not only to the alignment of feature points within the grid but also to the preservation of similarity transformations between sampled triangles. This approach, thus, succeeds in safeguarding the geometric structures within the images. However, GES-GSP can only protect individual structures and lacks the ability to comprehend the relationships between structures at a global level. As shown in Fig. 1 (c), it fails to preserve the overall geometric structure.
In order to globally preserve the structure of objects and prevent distortion, we propose to leverage Segment Anything Model (SAM) [25] to extract object-level geometric information, and generate triangle meshes within each object to ensure the overall preservation of object shapes during the stitching process. SAM is capable of automatically segmenting common scenes and extracting semantically rich, high-quality, holistic geometric structures, as shown in Fig. 1 (b). SAM treats objects as cohesive entities, transcending the segmentation of individual line segments. Furthermore, SAM is proficient at extracting object outlines, surpassing the capabilities of conventional straight or curved lines and allowing for a more nuanced understanding of the relationships between geometric structures.
@@
Our approach capitalizes on object-level geometric information. Drawing inspiration from rigid object deformations [19] in computer vision, we generate triangle meshes within each object. During the stitching process, these triangle meshes are collectively subjected to similarity transformations, effectively preserving the structure of the objects. As demonstrated in Fig. 1 (c) and (d), our method excels in maintaining the overall structure of objects, preventing distortion of prominent linear elements within the images, resulting in a more natural and faithful stitching outcome. Notably, even in scenarios where SAM fails to extract structures below the red dashed line, our approach remains effective by preserving the holistic structure of the object with triangle meshes.
Our proposed approach has resulted in a new state-of-the-art in the field of image stitching. The main contributions of this paper are as follows:
@@
  • We introduce semantic information into GES-GSP [13] through SAM [25], enabling the extraction of semantically enriched, uniformly distributed object-level image structures.
@@
  • We generate triangular meshes within objects and constrain them as a whole to undergo similarity transformations, preserving object shapes globally.
  • Through experiments on standard image stitching datasets, our method surpasses GES-GSP and establishes a new state-of-the-art benchmark.
2 Related works
Grid-based image stitching algorithm. Autostitch [3], a pioneering work in image stitching, identifies matching feature points and aligns them by homography transformation. Building upon this foundation, numerous stitching algorithms partition images into grids, compute geometric transformation relationships for each grid, and combine them into a global transformation to align overlapping regions and seamlessly transit the transformation to non-overlapping areas.
APAP [47] warps aim to maintain global projectivity while allowing local non-projective deviations to account for deviations from assumed imaging conditions. AANAP [28] linearizes the homography and gradually shifts it towards global similarity to mitigate perspective distortion. GSP [7] achieves more natural stitching results by leveraging local or global similarity transformations. However, grid deformation introduces some geometric distortion. To address this, LPC [20] proposed to leverage line-point consistency to preserve structures, introducing global collinear structures to enhance desirable characteristics for image warping. GES-GSP [13] not only preserves straight-line shapes but also protects the image’s curved structures, allowing curves to undergo similarity transformations to prevent distortion.
@@ 2 Related works
Nevertheless, LPC [20] and GES-GSP’s [13] extraction of straight or curved lines fails to fully capture the image’s geometric structure. Their structure preservation algorithms only protect individual lines without recognizing the overall structure. To address this, we employ semantic segmentation to extract meaningful geometric structures, use triangular grids to protect the overall object-level geometric structure, and establish connections between dispersed geometric transformations, achieving superior results.
@@
Geometric structure extraction. In LPC [20] and GES-GSP [13], the Line Segment Detector (LSD) [40] algorithm detects straight lines in images. GES-GSP employs edge detection methods like Canny [4] and HED [42] to identify edges, which represent the geometric structure of curves. LSD efficiently detects line segments based on gradient information. Canny is an image processing technique that identifies edges by detecting significant changes in pixel intensity, known for its noise suppression and precise edge mapping capabilities. HED is a neural network-based end-to-end edge detection method that outperforms Canny in edge detection. In recent years, deep learning-based semantic segmentation [38, 37, 6, 18, 46, 8] and instance segmentation [16, 17, 2, 23] methods have advanced rapidly, capable of segmenting objects with semantic information in images. However, they require large datasets [9, 31, 39] for training and can only segment fixed categories within the dataset. Thus, large-scale pretrained encoders [5, 35, 36] have been proposed under the encoder-decoder architecture, capable of fine-tuning for downstream tasks. SAM [25], introduced in early 2023, is a large model capable of semantic segmentation for point, crop box, and text prompts. SAM’s automatic mode automatically selects point prompts, segments images, and extracts object information.
We employ SAM [25] to extract object-level geometric structures and preserve these shapes during stitching. Compared to LSD [40], Canny [4], and HED [42], SAM’s extracted geometric structures carry semantic meaning, distribute evenly, and have a pleasing visual appearance. LPC [20] and GES-GSP [13] only protect independent, scattered geometric lines, while we protect objects composed of multiple geometric lines as a whole, resulting in superior performance.
3 GES-GSP pipeline
@@
GES-GSP [13] uses a grid mesh to guide the image deformation, where $V$ and $E$ represent the sets of vertices and edges within the grid mesh. GES-GSP aims to find a set of deformed vertex positions, denoted as $\widetilde{V}$, that minimizes the energy function $\psi(V)$. In summary, GES-GSP [13] builds upon the foundation of GSP [7] by introducing the GES term, specifically aimed at preserving the shape of lines, in addition to the Alignment term, local similarity term, and global similarity term within GSP [7]. In this section, we provide a brief overview of the optimization equations employed in GES-GSP [13].
@@
Alignment term extracts feature points from the input images with the goal of aligning these points after geometric transformation. For each feature point pair $(p,\Phi(p))$, $\widetilde{v}(p)$ represents the position of $p$ as a linear combination of four vertex positions, and $M$ represents the set of all feature point pairs. The algorithm linearly combines the coordinates of the four vertices of each grid to represent the position of $p$ through bilinear interpolation. By optimizing the positions of grid vertices after geometric transformation, it aims to bring $p$ as close as possible to $\Phi(p)$. Therefore, the energy equation is defined as:
@@
Local similarity term aims to ensure that the transition from overlapping to non-overlapping regions is natural and undistorted. Each grid undergoes a similarity transformation to minimize shape distortion. For an edge $(j,k)$, $S_{jk}$ represents its similarity transformation. Suppose $v_{j}$ transforms to $\widetilde{v_{j}}$ after deformation, and the energy function is defined as:
@@
For a detailed description of the four energy terms mentioned above, please refer to GSP and GES-GSP for more information. Ultimately, the sought-after vertices are determined by minimizing the following optimization equation for two images:
@@
4 Proposed method
In GES-GSP [13], it is necessary to extract large-scale geometric structures, including straight lines and curves, from the entire image and preserve these structures during stitching. Therefore, GES-GSP employs an edge detection algorithm to extract edges from the image. Subsequently, the image undergoes skeletonization, noise reduction, branch pruning, and corner point removal operations. Afterward, the LSD [40] algorithm is applied to detect straight lines within the image. Post-processing techniques are then employed to handle invalid edge information, including branch pruning, reconnection of discontinuous lines, and collinearity constraints. While this conventional approach can extract some geometric structures, its results often appear meaningless to human observers and fail to capture the complete geometric information of the image.
@@ 4 Proposed method
After obtaining the contours, we generate a triangular mesh for each semantic object, preserving the shape of the object through similarity transformations within the triangle mesh. Unlike the As-Rigid-As-Possible (ARAP) [19] method, we simplify computational complexity by directly locating the center of the object and connecting it to sampling points on the object’s semantic boundary to form a triangular mesh. In Fig 2, $V_{0}$ represents the object’s center, while $V_{1}$ and $V_{2}$ are sampling points on the semantic boundary of the object, forming a triangular mesh with these three points. One vertex, $V_{2}$, of the triangle can be represented using the edges of the triangle $\vec{V_{0}V_{1}}$ and an orthogonal coordinate system obtained by rotating this edge counterclockwise by 90 degrees:
@@
$\psi_{ges-o}(V)=\sum_{\beta=1}^{N_{c}}\sum_{\alpha=1}^{N_{s}}{\omega_{\alpha}^{\beta}E_{\alpha}^{\beta}}.$ (7)
@@
$\begin{split}\widetilde{V}&=\underset{\widetilde{V}}{arg\min}\left(\psi_{a}(\widetilde{V})+\lambda_{l}\psi_{l}(\widetilde{V})\right.\\ &\quad\left.+\psi_{g}(\widetilde{V})+\lambda_{ges-o}\psi_{ges-o}(\widetilde{V})\right).\end{split}$ (8)
For fair comparison, our parameters are identical to those of GES-GSP: $\lambda_{l}=0.75$, $\lambda_{ges-o}=1.5$. Our $\lambda_{ges-o}$ corresponds $\lambda_{ges}$ in to GES-GSP.
Models / Datasets AANAP [28] APAP [47] CAVE [22] DFW [27] DHW [14] GES-GSP [13]
Mean Distorted Residuals (MDR)
GSP [7] ECCV’16 1.06183 1.25495 0.90884 0.98457 1.08755 1.06986
GES-GSP [13] CVPR’23 1.06213 1.24249 0.90821 0.98034 1.05619 1.04473
OBJ-GSP (ours) 1.05930 1.20123 0.89731 0.97259 1.00496 0.98288
Improvement(%) 0.3 3.3 1.2 0.8 4.9 5.9
Naturalness Image Quality Evaluator (NIQE)
GSP 2.84241 3.4356 4.04708 5.61905 2.75485 3.84897
GES-GSP 2.77220 3.48713 4.03835 5.71544 2.70838 3.79240
OBJ-GSP (ours) 2.74965 3.39280 4.01565 5.69104 2.60825 3.70041
Improvement(%) 0.8 2.7 0.6 0.4 3.7 2.4
Models / Datasets LPC [20] REW [26] SEAGULL [30] SVA [32] SPHP [29] Our dataset
Mean Distorted Residuals (MDR)
GSP 1.30562 1.16192 1.14467 1.51158 1.17784 1.15296
GES-GSP 1.22256 1.15462 1.11993 1.47197 1.14340 1.14366
OBJ-GSP (ours) 1.10622 1.08635 1.08296 1.47813 1.07699 1.12229
Improvement(%) 9.5 5.9 3.3 -0.4 5.8 1.9
Naturalness Image Quality Evaluator (NIQE)
GSP 4.28546 3.18549 5.10784 7.00495 3.04781 2.66597
GES-GSP 3.35315 2.75234 4.69390 6.96670 2.97173 2.64986
OBJ-GSP (ours) 3.23057 2.81480 4.08903 6.96149 2.49712 2.54906
Improvement(%) 3.7 -2.2 12.9 0.5 16.0 3.8
Table 1: Quantitative results on 9 image stitching datasets. We report the mean MDR and NIQE on each dataset. We label best results with bold text. A lower MDR and NIQE indicates better stitched panorama. Improvement row compares our proposed method and GES-GSP. The whole dataset consists of 124 images and our mean improvement to GES-GSP is 3.5% in MDR and 3.8% in NIQE.

### v2

@@ Abstract
The topic of stitching images with globally natural structures holds paramount significance, with two main goals: alignment and distortion prevention. The existing approaches exhibit the ability to align well, yet fall short in maintaining object structures. In this paper, we endeavour to safeguard the overall OBJect-level structures within images based on Global Similarity Prior (OBJ-GSP), while concurrently ensuring a good alignment performance. Our approach leverages the segment anything model [28] to extract geometric structures with semantic information. We employ triangular meshes in image transformation to protect the overall shapes of objects within images. The experimental results obtained on standard benchmarks demonstrate that OBJ-GSP outperforms existing methods in both alignment and shape preservation. Our implementation and dataset are publicly available at https://github.com/RussRobin/OBJ-GSP.
@@ 1 Introduction
Image stitching aims to align multiple images and create a composite image with a larger field of view. This method is widely utilized across diverse domains, including smartphone panoramic photography [52], robotic navigation [11], and virtual reality [1, 27]. The image stitching research community grapples with two fundamental challenges:
  • Alignment: matching corresponding points in multiple images so that they are seamlessly blended together in a panoramic image.
  • Distortion prevention: preventing structures in original images from being distorted in the panorama.
In recent years, the problem of alignment has largely been addressed [40]. Methods such as APAP [58] and GSP [7] divide the images into multiple grids, compute local transformation matrices within each grid, and combine them with global transformation matrices to achieve precise alignment in overlapping regions. However, they tend to pursue the alignment of feature points excessively without recognizing the shapes of objects in the images, thus distorting the shapes and leading to unnatural stitching results, as depicted in Fig. 1 (a)-(c).
Figure 1: Red boxes indicate blurriness. (a) and (b) are not aligned well. (c) GSP [7] aligns well but distorts the building. Based on this, (d) GES-GSP [14] tries to prevent distortion but still fails in this case. (f) our method protects the structure of the building by sampling on object contours extracted by SAM (e).
The main concern of image stitching is to prevent distortion on the basis of good alignment performance [14, 24, 40]. Our key insight is to globally extract the structure of objects in the input images and preserve them during stitching.
@@
Specifically, we extract object structures by SAM [28] and preserve them by optimizing the deformation of triangle meshes. Drawing inspiration from rigid object deformations [22] in computer vision, we generate triangle meshes within each object. During image transformation, these triangle meshes tend to reach a balance between projection and similarity transformations, effectively preserving the structure of the objects. As demonstrated in Fig. 1 (f), our method excels in maintaining the overall structure of objects, preventing distortion of prominent shapes within the original images.
Our approach capitalizes on object-level geometric information, and SAM extracts object-level geometric information. As shown in Fig. 1 (e), SAM treats objects as cohesive entities, transcending the segmentation of individual lines and curves adopted in previous works [14, 24]. This allows for a more nuanced understanding of the relationships between individual geometric structures.
To summarize, the main contributions of the proposed OBJ-GSP method include:
  • We propose to preserve object shapes during image transformation to stitch a natural panorama. Our method reaches a balance between projection (to align) and similarity (to prevent distortion) transformation.
@@
  • We introduce semantic information into image stitching through SAM [28], enabling the extraction of semantically enriched, uniformly distributed object-level image structures.
@@
  • We generate triangular meshes within objects and constrain them as a whole to undergo similarity transformations, preserving object shapes globally. Through experiments on standard image stitching datasets, our method surpasses existing methods in both alignment and shape preservation.
2 Related work
Grid-based image stitching. Autostitch [3], a pioneering work in image stitching, matches feature points and aligns them by homography transformation. Building upon this foundation, numerous stitching algorithms partition images into grids, compute geometric transformation relationships for each grid, and combine them into a global transformation to align overlapping regions and seamlessly transit the transformation to non-overlapping areas. APAP [58], AANAP [31], and GSP [7] have evolved over time, essentially addressing most alignment problems in images. However, their grid deformation methods have no knowledge of object shapes. They pay too much attention on alignment and thus causes geometric distortion. To address this, LPC [24] proposes to leverage line-point consistency to preserve structures, introducing global collinear structures to enhance desirable characteristics for image warping. GES-GSP [14] not only preserves straight-line shapes but also protects the image’s curved structures, allowing curves to undergo similarity transformations to prevent distortion.
@@ 2 Related work
Nevertheless, LPC [24] and GES-GSP’s [14] extraction of straight or curved lines fails to fully capture the image’s geometric structure. Their structure preservation algorithms only protect individual lines without recognizing the overall structure. To address this, we employ semantic segmentation to extract meaningful geometric structures. We use triangular grids to protect the overall object-level geometric structure, and establish connections between dispersed geometric transformations, achieving superior results.
@@
Geometric structure extraction. In LPC [24] and GES-GSP [14], the Line Segment Detector (LSD) [17] algorithm detects straight lines in images. GES-GSP employs edge detection methods like Canny [4] and HED [51] to identify edges, which represent the geometric structure of curves. In recent years, deep learning-based semantic segmentation [45, 43, 6, 21, 57, 8] and instance segmentation [18, 20, 2, 26] methods have advanced rapidly, capable of segmenting objects with semantic information in images. However, they require large datasets [10, 34, 47] for training and can only segment fixed categories within the dataset. Thus, large-scale pretrained encoders [5, 41, 42] have been proposed under the encoder-decoder architecture, capable of fine-tuning for downstream tasks. SAM [28], introduced in early 2023, is a large model capable of semantic segmentation for point, crop box, and text prompts. SAM’s automatic mode automatically selects point prompts, segments images, and extracts object information. We employ SAM [28] to extract object-level geometric structures and preserve these shapes during stitching. Compared to LSD [17], Canny [4], and HED [51], SAM’s extracted geometric structures carry semantic meaning, distribute evenly, and have a pleasing visual appearance. LPC [24] and GES-GSP [14] only protect independent, scattered geometric lines, while we protect objects composed of multiple geometric lines as a whole, resulting in superior performance. In the future, the accuracy and speed of SAM-type methods will both improve [53, 59], further enhancing the quality and speed of our image stitching techniques.
Deep-learning based stitching and OBJ-GSP. In recent years, several methods [23, 40] like UDIS [39] have attempted to model certain image stitching steps as unsupervised deep learning problems, leading to notable advances in this field. UDIS++ [40] also addresses the distortion problem on the basis of good alignment performance, which aligns perfectly with our goals. It attempts to replace feature points extracted by traditional methods with deep learning-based backbones to reconstruct panoramas through feature map reconstruction, and they specifically design distortion-solving losses based on grid transformation. We adhere to the traditional approach in the stitching domain by preserving results in grid transformation, while UDIS++ provides a completely new deep learning-based pipeline, although currently its performance is not as good as ours. We provide a detailed comparison with UDIS++ in the supplementary material.
3 A revisit of stitching pipeline
In this section, we revisit the previous state-of-the-art stitching pipeline by GES-GSP [14]. In the next section we will show how OBJ-GSP preserves object-level structures in this pipeline.
@@ 3 A revisit of stitching pipeline
GES-GSP uses a grid mesh to guide the image deformation, where $V$ and $E$ represent the sets of vertices and edges within the grid mesh. GES-GSP aims to find a set of deformed vertex positions, denoted as $\widetilde{V}$, that minimizes the energy function $\psi(V)$. In short, GES-GSP [14] builds upon the foundation of GSP [7] by introducing the GES term, specifically aimed at preserving the shape of lines, in addition to the Alignment term, Local similarity term, and Global similarity term within GSP [7]. In this section, we provide a brief overview of the optimization equations employed in GES-GSP [14].
@@
Alignment term extracts feature points from the input images with the goal of aligning these points after geometric transformation. For each feature point pair $(p,\Phi(p))$, $\widetilde{v}(p)$ represents the position of $p$ as a linear combination of four vertex positions, and $M$ represents the set of all feature point pairs. The algorithm linearly combines the coordinates of the four vertices of each grid to represent the position of $p$ through bi-linear interpolation. By optimizing the positions of grid vertices after geometric transformation, it aims to bring $p$ as close as possible to $\Phi(p)$. Therefore, the energy equation is defined as:
@@
Local similarity term aims to ensure that the transition from overlapping to non-overlapping regions is natural. Each grid undergoes a similarity transformation to minimize shape distortion. For an edge $(j,k)$, $S_{jk}$ represents its similarity transformation. Suppose $v_{j}$ transforms to $\widetilde{v_{j}}$ after deformation, and the energy function is defined as:
@@
For a detailed description of the four energy terms mentioned above, please refer to GSP [7] and GES-GSP [14] for more information. Ultimately, the sought-after vertices are determined by minimizing the following optimization equation for two images:
@@
4 The proposed method
In GES-GSP [14], it is necessary to extract large-scale geometric structures, including straight lines and curves, from the entire image and preserve these structures during stitching. Therefore, GES-GSP employs an edge detection algorithm to extract edges from the image. Subsequently, the image undergoes skeletonization, noise reduction, branch pruning, and corner point removal operations. Afterward, the LSD [17] algorithm is applied to detect straight lines within the image. Post-processing techniques are then employed to handle invalid edge information, including branch pruning, re-connection of discontinuous lines, and collinearity constraints. While this conventional approach can extract some geometric structures, its results often appear meaningless to human observers and fail to capture the complete geometric information of the image.
@@ 4 The proposed method
After obtaining the contours, we generate a triangular mesh for each semantic object, preserving the shape of the object through similarity transformations within the triangle mesh. Unlike the As-Rigid-As-Possible (ARAP) [22] method, we simplify computational complexity by directly locating the center of the object and connecting it to sampling points on the object’s semantic boundary to form a triangular mesh. In Fig 2, $V_{0}$ represents the object’s center, while $V_{1}$ and $V_{2}$ are sampling points on the semantic boundary of the object, forming a triangular mesh with these three points. $(x_{01},y_{01})$ refer to the known coordinates of a feature point in the local coordinate plane. One vertex, $V_{2}$, of the triangle can be represented using the edges of the triangle $\vec{V_{0}V_{1}}$ and an orthogonal coordinate system obtained by rotating this edge counterclockwise by 90 degrees:
@@
$\psi_{obj}(V)=\sum_{\beta=1}^{N_{c}}\sum_{\alpha=1}^{N_{s}}{\omega_{\alpha}^{\beta}E_{\alpha}^{\beta}}.$ (7)
@@
$\begin{split}\widetilde{V}&=\underset{\widetilde{V}}{arg\min}\left(\psi_{a}(\widetilde{V})+\lambda_{l}\psi_{l}(\widetilde{V})\right.\\ &\quad\left.+\psi_{g}(\widetilde{V})+\lambda_{obj}\psi_{obj}(\widetilde{V})\right).\end{split}$ (8)
For fair comparison, our parameters are identical to those of GES-GSP: $\lambda_{l}=0.75$, $\lambda_{obj}=1.5$. Our $\lambda_{obj}$ corresponds $\lambda_{ges}$ in to GES-GSP.
Table 1: We report the mean MDR to measure alignment performance, and NIQE for distortion prevention evaluation. UDIS and UDIS++ are not feature point based so we only report NIQE and leave qualitative results in supplementary material. Best results are labeled with bold text. Lower MDR and NIQE indicates better stitched panorama. Improvement row compares our proposed method and GES-GSP. Our mean improvement to GES-GSP is 3.5% in MDR and 3.8% in NIQE.
Models / Datasets Our dataset AANAP [31] APAP [58] CAVE [25]
Mean Distorted Residuals (MDR $\downarrow$)
GSP [7] ECCV16 1.15296 1.06183 1.25495 0.90884
GES-GSP [14] CVPR22 1.14366 1.06213 1.24249 0.90821
OBJ-GSP (ours) 1.12229 1.05930 1.20123 0.89731
Improvement(%) 1.9 0.3 3.3 1.2
Naturalness Image Quality Evaluator (NIQE $\downarrow$)
UDIS [39] TIP21 3.69421 3.01517 3.69421 -
UDIS++ [40] ICCV23 3.34003 2.95493 3.56812 4.07702
GSP 2.66597 2.84241 3.4356 4.04708
GES-GSP 2.64986 2.77220 3.48713 4.03835
OBJ-GSP (ours) 2.54906 2.74965 3.39280 4.01565
Improvement(%) 3.8 0.8 2.7 0.6
Models / Datasets DFW [30] DHW [15] GES-GSP [14] LPC [24]
Mean Distorted Residuals (MDR $\downarrow$)
GSP 0.98457 1.08755 1.06986 1.30562
GES-GSP 0.98034 1.05619 1.04473 1.22256
OBJ-GSP (ours) 0.97259 1.00496 0.98288 1.10622
Improvement(%) 0.8 4.9 5.9 9.5
Naturalness Image Quality Evaluator (NIQE $\downarrow$)
UDIS 5.74137 3.28645 5.02442 3.76994
UDIS++ 5.09680 3.23392 4.93279 3.66565
GSP 5.61905 2.75485 3.84897 4.28546
GES-GSP 5.71544 2.70838 3.79240 3.35315
OBJ-GSP (ours) 5.69104 2.60825 3.70041 3.23057
Improvement(%) 0.4 3.7 2.4 3.7
Models / Datasets REW [29] SEAGULL [33] SVA [35] SPHP [32]
Mean Distorted Residuals (MDR $\downarrow$)
GSP 1.16192 1.14467 1.51158 1.17784
GES-GSP 1.15462 1.11993 1.47197 1.14340
OBJ-GSP (ours) 1.08635 1.08296 1.47813 1.07699
Improvement(%) 5.9 3.3 -0.4 5.8
Naturalness Image Quality Evaluator (NIQE $\downarrow$)
UDIS 3.61888 4.67437 8.02090 4.35149
UDIS++ 3.67661 4.38520 7.5419 4.1248
GSP 3.18549 5.10784 7.00495 3.04781
GES-GSP 2.75234 4.69390 6.96670 2.97173
OBJ-GSP (ours) 2.81480 4.08903 6.96149 2.49712
Improvement(%) -2.2 12.9 0.5 16.0